solve the system of equations using the substution method
-5x-y=18
x+2=9
solution(?,?)

Answer by rfer(16322)

x=7
-5x-y=18
-5(7)-y=18
-35-y=18
y=-53
------------
(7,-53)
should the problem be x+2y=9??